# Faster Cascades via Speculative Decoding

### v1

@@ 3.2 Optimal token-level cascade deferral
While Chow’s rule (2) is easy to implement, it can be sub-optimal if the smaller model’s maximum token probability is not reflective of which of the two models are better equipped to predict the next token for a given prefix [22]. Given this, it is natural to ask what the optimal deferral rule $r$ for a token-cascade looks like, and whether we can reasonably approximate this rule.
@@
For this, we must first specify an objective to minimize at each step $t$. Following the prior cascade literature [22, 18], a reasonable objective to minimize is the expected loss from the deferral rule against the ground-truth distribution $\mathbb{P}$, with an added cost for deferring to the larger model. We state this below for a fixed prefix $x_{<t}$, using the short-hand $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$ and $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$:
$L_{\rm def}(r;x_{<t})$ $=\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\big{(}\ell(x_{t},p_{t})+\alpha\big{)}\Big{]},$ (3)
@@
for a cost penalty $\alpha\geq 0$ and loss function $\ell:\mathscr{V}\times\Delta_{\mathscr{V}}\rightarrow\mathbb{R}_{+}$. Common choices for $\ell$ include the 0-1 loss $\ell_{\text{0-1}}(x_{t},q_{t})=\mathbf{1}\left(x_{t}\neq\operatorname{argmax}_{v}q_{t}(v)\right)$ and the log loss $\ell_{\log}(x_{t},q_{t})=-\log\left(q_{t}(x_{t})\right).$
Lemma 1 (Optimal deferral for token-level cascades [22]).
The minimizer of (3) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha.$ (4)
@@
Intuitively, we compare the expected loss from $q$ with the expected cost of invoking $p$, and decide to defer when the latter is smaller. We note here that this optimization problem is set up for a fixed prefix $x_{<t}$. One may also consider the coupled optimization problem across all prefixes from 1 to $T$.
@@
Plug-in estimator for (4). The optimal rule in (4) requires computing expectations over the ground-truth distribution $\mathbb{P}(\cdot|x_{>t})$, which is not available during inference time. A common approach in the cascades literature is to replace the expected losses with the models’ confidence estimates [22]. For example, when $\ell=\ell_{\text{0-1}}$, it may be reasonable to use $1-\max_{v}q_{t}(v)$ as an estimate of the expected 0-1 loss $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$ and $1-\max_{v}p_{t}(v)$ as an estimate of $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$. The extent to which these estimates are accurate depend on how well $q$ and $p$ are calibrated [17]. The resulting plug-in estimator for (4) is given by:
$\textstyle\hat{r}_{\rm\tt Diff}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha.$ (5)
@@
Similarly, when $\ell=\ell_{\log}$, we may use the entropy $-\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))$ from $q_{t}$ as an estimate of its expected log-loss, and similarly for $p_{t}$ (see §C).
@@
$\hat{r}_{\rm\tt Diff}$ cannot be directly used in a token-level cascade, as it needs the larger model to be invoked at every position $t$. However, it serves as an oracle that allows to analyze the head-room available to improve upon Chow’s rule. See also Remark 2.
3.3 When do token-level cascades outperform speculative decoding?
@@
Token-level cascades and speculative decoding differ in the distribution over tokens they seek to mimic. Speculative decoding seeks to mimic the larger model’s output distribution (or an approximation to it). As a result, the quality of the sampled output is limited by the accuracy of the larger model. On the other hand, token-level cascades seek to output distributions that closely approximate the ground-truth label distribution, and can potentially yield better quality than even the larger model.
@@
Indeed, speculative decoding would be ideal in applications where the verification model $p$ is uniformly better than the draft model $q$ on all inputs. However, when there are inputs where the draft model fares better than the verifier, one may want to retain the drafter’s predictions even when it disagrees with the verifier. Similarly, in settings where both the drafter and verifier fare poorly on some inputs (e.g., due to label noise), one may again want to ignore the disagreement between the drafter and verifier, and avoid triggering unnecessary roll-backs.
@@
As a concrete example, we consider token-level cascades of T5 models [35] of two different sizes finetuned on a WMT EN $\rightarrow$ DE translation [3] and an extreme summarization (XSum) task [31]. We construct these cascades using both (token-level) Chow’s rule in (2) and the oracle Diff rule in (5), and also apply speculative decoding with the smaller (larger) model as the drafter (verifier). In Figure 1, we plot quality as a function of fraction of samples deferred to the large model, as we vary the cost parameter $\alpha$. Note that with speculative decoding, each verification step is counted as a single call to the large model. While speculative decoding matches the quality of the large model (right-most point), the oracle rule yields significantly better quality on a wide range of operating points. Even Chow’s rule, which is sub-optimal for cascading [22], outperforms speculative decoding in a small region.
@@
However, as also evident from the plots, token-level cascades require a significantly larger number of calls to the larger model to achieve the same quality. This is because token-level cascades are executed sequentially: whenever $q$ defers, we execute $p$ once to generate one next token for the prefix accumulated so far, and the control transfers back to $q$. In contrast, speculative decoding runs $p$ in scoring mode to verify multiple draft tokens from $q$ in parallel. Moreover, the stochastic verification algorithm in speculative decoding often results in fewer tokens from $q$ getting rejected compared to the deterministic deferral rules used in a cascade. These observations motivate a natural question: given their complementary strengths, how can we leverage the best of both these techniques?
@@
Figure 1: Plot of quality as a function of deferral rate for cascades constructed from T5 models (under temperature sampling with $T=1$). The left-most point represents the small model and the right-most represents the large model. We compare token-level cascades constructed with Chow’s rule (Chow) and an oracle deferral rule (Diff), and speculative decoding with block size $\gamma=5$. We measure deferral rate in terms of the number of calls to the larger model; with a cascade, each call yields exactly one token, whereas with speculative decoding, a single call scores $\gamma$ draft tokens in parallel. While speculative decoding matches the quality of the large model (see dashed horizontal line), the oracle deferral rule yields significantly better quality on a range of deferral rates; this however comes at the cost of higher number of calls to the large model.
@@ 4 Speculative Cascades: Leveraging the Best of Both Worlds
In addressing the above question, we present our main contribution: a principled approach to combining the superior quality of cascades with the faster execution of speculative decoding approaches.
@@ 4.1 Speculative decoding with general target distributions
This general procedure not only encompasses standard speculative decoding [26] for $\mathbb{T}(q,p)=p$, but also includes lossy speculative decoding [44] as a special case:
@@
Algorithm 4 reduces to the lossy speculative sampling procedure in [44] with parameters $\alpha$ and $\beta$ when $\mathbb{T}(q,p)(v)=\max\{\min\{q(v),\frac{p(v)}{1-\alpha}\},\frac{p(v)}{\beta}\}$.
@@ 4.2 From sequential to speculative cascades
Equipped with Algorithm 4, we now propose new cascading techniques that implement their deferral rule in a speculative manner. Recall that a token-level cascade of two models $q$ and $p$ is defined by a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$. For a prefix $x_{<t}$, the next-token distribution at position $t$ modeled by this cascade can be written as:
@@
for some function $\delta:\Delta_{\mathscr{V}}\times\Delta_{\mathscr{V}}\rightarrow\{0,1\}$ that maps distributions $(q,p)$ to a binary decision. For example, for Chow, $\delta(q,p)=\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$, and for Diff, $\delta(q,p)=\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}.$ See Table 1 for a summary of target distributions for different deferral rules.
@@
Unlike a sequential cascade, where the larger model’s distribution $p$ is not available at the time the deferral decision is made (see Remark 1), with a speculative cascade, we can accommodate deferral rules like Diff that depend on both $q$ and $p$. This is because we run the larger model $p$ in parallel on drafts generated by the smaller model $q$, allowing us to compute both $p(\cdot)$ and $q(\cdot)$ on every prefix.
@@
    $q_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}q(\cdot|x_{<t+j});~{}~{}~{}~{}$ $x_{t+j}\sim q_{t+j}(\cdot)$
@@
$p_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}p(\cdot|x_{<t+j}),~{}\forall j\in[\gamma]$
@@
$a_{j}\sim\text{Ber}\left(\min\left\{1,\frac{\pi_{t+j}({x_{t+j}})}{q_{t+j}({x_{t+j}})}\right\}\right),~{}\forall j\in[\gamma-1];$    $a_{\gamma}=0$
@@
$p_{\rm res}(\cdot)=\begin{cases}\operatorname{norm}\left(\max\left\{0,\,\pi_{t+j^{*}}(\cdot)-q_{t+j^{*}}(\cdot)\right\}\right)&\text{if}~{}j^{*}<\gamma\\ p_{t+\gamma}(\cdot)&\text{else}\end{cases}$
@@
4.3 Deferral risk for speculative cascades
As with sequential cascades (§2), we begin by defining an objective to minimize. We seek a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that minimizes a loss against the ground-truth distribution, while limiting the inference cost to be within a budget. (Per above, this deferral rule implicitly defines a target distribution $\pi$.) The inference cost crucially depends on how frequently a draft token is rejected in the verification phase, triggering a rollback. To this end, we derive the probability that a token sampled from $q$ is rejected during verification, for a target distribution resulting from a deferral rule $r$.
@@ 4.3 Deferral risk for speculative cascades
For a given prefix $x_{<t}$, and target distribution $\pi=(1-r(x_{<t}))\cdot q_{t}+r(x_{<t})\cdot p_{t}$, the probability of a token drawn from draft distribution $q_{t}$ being rejected is equal to:
$r(x_{<t})\cdot D_{\textup{{TV}}}(p_{t},q_{t}),$
where $D_{\textup{{TV}}}(p,q)=\sum_{v\in\mathscr{V}}\max\{0,p(v)-q(v)\}$ is the TV distance between $p$ and $q$.
Intuitively, whenever $r(x_{<t})=0$, $\pi(v)=q_{t}(v)$ and therefore there is no rejection or roll-back; when $r(x_{<t})=1$, the rejection rate equals $D_{\textup{{TV}}}(p_{t},q_{t})$.
@@
$\min_{r}$ $~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\ell(x_{t},p_{t})\big{)}\Big{]}$
[…]
for some budget $B>0$. Equivalently, one may minimize an unconstrained objective similar to (3), for suitable cost parameter $\alpha>0$ (see §C.4):
$L_{\rm spec}(r;x_{<t})$ $=\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\big{(}\ell(x_{t},p_{t})+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})\big{)}\Big{]},$ (8)
Contrasting (8) with the deferral risk in (3) for a sequential cascade, a key difference is that the cost of deferring to the larger model is no longer a constant, but depends on the similarity between $q_{t}$ and $p_{t}$, as measured by the TV distance between them.
4.4 Optimal speculative deferral
We next derive the optimal deferral rule for (8), and construct a feasible estimator for it.
@@ 4.4 Optimal speculative deferral
The minimizer of (8) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (9)
@@
Plug-in estimator for (9). The optimal rule requires estimating expectations with respect the ground-truth distribution $\mathbb{P}(\cdot|x_{<t}).$ We employ similar plug-in estimators as the ones used with sequential cascades (§3). When $\ell=\ell_{\text{0-1}}$, we replace the expected 0-1 loss with (one minus) the maximum probability from the model, giving us:
$\textstyle\hat{r}_{\rm\tt OPT}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (10)
@@
$L_{\rm spec}(\hat{r}_{\rm\tt OPT};x_{<t})-\min_{r}\,L_{\rm spec}(r;x_{<t})~{}\leq~{}\max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,\max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$

### v2

@@ 3.2 Optimal token-level cascade deferral
While Chow’s rule equation 2 is easy to implement, it can be sub-optimal if the smaller model’s max-token probability is not reflective of which of the two models are better equipped to predict the next token for a given prefix (Jitkrittum et al., 2023). Given this, it is natural to ask what the optimal deferral rule $r$ for a token-cascade looks like, and whether we can reasonably approximate this rule.
@@
For this, we must first specify an objective to minimize at each step $t$. Following the prior cascade literature (Jitkrittum et al., 2023; Gupta et al., 2024), a reasonable objective to minimize is the expected loss from the deferral rule against the ground-truth distribution $\mathbb{P}$, with an added cost for deferring to the larger model. We state this below for a fixed prefix $x_{<t}$, using as before color=RoyalBlue!35!white]“using once more”? Hari: done! the short-hand $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$ and $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$:
$L_{\rm def}(r;x_{<t})$ $=\mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(v,q_{t})+r(x_{<t})\cdot\big{(}\ell(v,p_{t})+\alpha\big{)}\Big{]},$ (3)
@@
for a cost penalty $\alpha\geq 0$ and loss function $\ell:\mathscr{V}\times\Delta_{\mathscr{V}}\rightarrow\mathbb{R}_{+}$. Common choices for $\ell$ include the 0-1 loss $\ell_{\text{0-1}}(v,q_{t})=\bm{1}\left(v\neq\operatorname*{arg\,max}_{v^{\prime}}q_{t}(v^{\prime})\right)$ and the log loss $\ell_{\log}(v,q_{t})=-\log\left(q_{t}(v)\right).$
Lemma 1 (Optimal deferral for token-level cascades (Jitkrittum et al., 2023)).
The minimizer of equation 3 is of the form:
$r^{*}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(v,q_{t})\right]\,>\,\mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(v,p_{t})\right]+\alpha.$ (4)
@@
Intuitively, we compare the expected loss from $q$ with the expected cost of invoking $p$, and decide to defer when the latter is smaller. We note here that this optimization problem is set up for a fixed prefix $x_{<t}$. One may also consider the coupled optimization problem across all positions.
@@
Plug-in estimator for equation 4. The optimal rule in equation 4 requires computing expectations over the ground-truth distribution $\mathbb{P}(\cdot|x_{>t})$, which is not available during inference time. A common approach in the cascades literature is to replace the expected losses with the models’ confidence estimates (Jitkrittum et al., 2023). For example, when $\ell=\ell_{\text{0-1}}$, it may be reasonable to use $1-\max_{v}q_{t}(v)$ as an estimate of the expected 0-1 loss $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$ and $1-\max_{v}p_{t}(v)$ as an estimate of $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$. The extent to which these estimates are accurate depend on how well $q$ and $p$ are calibrated (Guo et al., 2017). The resulting plug-in estimator for (4) thresholds the difference of confidence estimates from both distributions:
$\textstyle{\hat{r}_{\rm\tt Diff}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha.}$ (5)
@@
Similarly, when $\ell=\ell_{\log}$, we may use the entropy $-\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))$ from $q_{t}$ as an estimate of its expected log-loss, and similarly for $p_{t}$ (see Appendix C).
@@
For efficiency reasons, $\hat{r}_{\rm\tt Diff}$ cannot be directly used in a token-level cascade,color=green!30!white]“to improve efficiency”? We can if we want quality, right? Hari: yep, added a clarification as it needs the large model to be invoked at every step $t$. However, it serves as an oracle that allows to analyze the head-room available to improve upon Chow’s rule. See also Remark 2.
3.3 Contrasting token-level cascade and speculative decoding trade-offs
@@
Token-level cascades and speculative decoding differ in the distribution over tokens they seek to mimic. Speculative decoding seeks to mimic the large model’s output distribution, and is usually used when one wants to match the quality of the large model. On the other hand, token-level cascades seek to output distributions that closely approximate the ground-truth label distribution and potentially offer good cost-quality trade-offs, sometimes yielding better quality than even the large model.
@@
Cascades are useful when the draft model fares better than the verifier on some inputs, and one may want to retain the drafter’s predictions even when it disagrees with the verifier. Even in cases where both the drafter and verifier fare poorly on some inputs (e.g., due to label noise), one may want to ignore the disagreement between the drafter and verifier to avoid triggering unnecessary roll-backs.
@@
As a concrete example, we consider token-level cascades of T5 models (Raffel et al., 2020b) of two different sizes finetuned on a WMT EN $\rightarrow$ DE translation Bojar et al. (2014b) and an extreme summarization (XSum) task (Narayan et al., 2018). We construct these cascades using both (token-level) Chow’s rule in equation 2 and the oracle Diff rule in equation 5, and also apply speculative decoding with the smaller (larger) model as the drafter (verifier). In Figure 1, we plot quality as a function of fraction of samples deferred to the large model (number of deferrals divided by number of generated tokens), as we vary the cost parameter $\alpha$. Note that with speculative decoding, each verification stepcolor=green!30!white]Minor: worth explicitly mentioning $\gamma$ tokens are verified in parallel. verifies $\gamma$ tokens in parallel, but is counted as a single deferral to the large model. While speculative decoding matches the quality of the large model (right-most point), the oracle yields significantly better cost-qualty trade-offs. Even Chow’s rule, which is sub-optimal for cascading (Jitkrittum et al., 2023), offers better trade-offs, and outperforms speculative decoding in a small region.color=cyan]But does this small improvement comes at some added cost? E.g., higher latency? – Okay, we seem to be discussing this in the following program. But should we breifly allude to this in the caption of Figure 1 as well? Hari: done! As noted by Kim et al. (2023), this may be attributed to the ensembling effect in a cascade. color=RoyalBlue!35!white]the caption refers to “deferral rate”, but the x-axis says “Fraction of calls”. Hari: changed this as per next comment! color=RoyalBlue!35!white]maybe the x-axis could be more explicit, like “# large-model calls / # generated tokens”. Hari: done!
@@
However, as also evident from the plots, token-level cascades require a significantly larger number of deferrals to the large model to achieve the same quality. This is because token-level cascades are executed sequentially: whenever $q$ defers, we execute $p$ once to generate one next token for the prefix accumulated so far, and the control transfers back to $q$. In contrast, speculative decoding runs $p$ in scoring mode to verify $\gamma$ draft tokens from $q$ in parallel. Moreover, the stochastic verification algorithm in speculative decoding often results in fewer tokens from $q$ getting rejected compared to the deterministic deferral rules used in a cascade. These observations motivate a natural question: given their complementary strengths, how can we leverage the best of both these techniques?
@@
Figure 1: Plots of quality as a function of the number of deferrals to the larger model divided by the total number of generated tokens for cascades constructed from T5 models (under temperature sampling with $T=1$). The left-most point represents the small model and the right-most represents the large model. We compare token-level cascades constructed with Chow’s rule (Chow) and an oracle deferral rule (Diff), and speculative decoding with block size $\gamma=5$. With a cascade, each call to the large model yields exactly one token, whereas with speculative decoding, a single call scores $\gamma$ draft tokens in parallel. While speculative decoding matches the quality of the large model (see dashed horizontal line), the oracle deferral rule yields significantly better quality on a range of deferral rates; this however comes at the cost of higher number of deferrals to the large model.
@@ 4 Speculative Cascades: Leveraging the Best of Both Worlds
In addressing the above question, we present our main contribution: a principled approach to combining the better trade-offs cascades offer with the faster execution of speculative decoding.
@@ 4.1 Speculative decoding with general target distributions
This general procedure not only encompasses standard speculative decoding (Leviathan et al., 2023) for $\mathbb{T}(q,p)=p$, but also includes lossy speculative decoding (Tran-Thien, 2023) as a special case:
@@
Algorithm 4 reduces to the lossy speculative sampling procedure in (Tran-Thien, 2023) with parameters $\alpha$ and $\beta$ when $\mathbb{T}(q,p)(v)=\max\{\min\{q(v),\frac{p(v)}{1-\alpha}\},\frac{p(v)}{\beta}\}$.
@@ 4.2 From sequential to speculative cascades
Equipped with Algorithm 4, we now propose new cascading techniques that implement their deferral rule in a speculative manner. Recall from §3.1 that a token-level cascade of two models $q$ and $p$ is defined by a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$. For a prefix $x_{<t}$, the next-token distribution at position $t$ modeled by this cascade can be written as:
@@
for some function $\delta:\Delta_{\mathscr{V}}\times\Delta_{\mathscr{V}}\rightarrow\{0,1\}$ that maps distributions $(q,p)$ to a binary decision. For example, for Chow, $\delta(q,p)=\bm{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$, and for Diff, $\delta(q,p)=\bm{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}.$ See Table 1 for a summary of target distributions for different deferral rules. color=green!30!white]Table 1 is referred to for the first time here. But it is placed a few pages back. Hari: added a ref in the contributions list
@@
In a sequential cascade, the large model’s distribution $p$ cannot be used color=green!30!white]Saying that $p$ is not available may confuse some readers. “using $p$ defeats the purpose of a cascade”? Hari: done! at the time the deferral decision is made (see Remark 1), as this would defeat the purpose of the cascade. With a speculative cascade, however, we can employ rules like Diff that depend on both $q$ and $p$. This is because we run the large model $p$ in parallel on drafts generated by the small model $q$, allowing us to compute both $p(\cdot)$ and $q(\cdot)$ on every prefix.
@@
    $q_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}q(\cdot|x_{<t+j});\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $x_{t+j}\sim q_{t+j}(\cdot)$
@@
$p_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}p(\cdot|x_{<t+j}),\leavevmode\nobreak\ \forall j\in[\gamma]$
@@
$a_{j}\sim\text{Ber}\left(\min\left\{1,\frac{\pi_{t+j}({x_{t+j}})}{q_{t+j}({x_{t+j}})}\right\}\right),\leavevmode\nobreak\ \forall j\in[\gamma-1];$    $a_{\gamma}=0$
@@
$p_{\rm res}(\cdot)=\begin{cases}\operatorname{norm}\left(\max\left\{0,\,\pi_{t+j^{*}}(\cdot)-q_{t+j^{*}}(\cdot)\right\}\right)&\text{if}\leavevmode\nobreak\ j^{*}<\gamma\\ \pi_{t+\gamma}(\cdot)&\text{else}\end{cases}$
@@
4.3 Optimal speculative cascade deferral
As with sequential cascades (§2), we begin by defining an objective to minimize. We seek a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that minimizes a loss against the ground-truth distribution, while limiting the inference cost to be within a budget. (Per above, this deferral rule implicitly defines a target distribution $\pi$.) The inference cost crucially depends on how frequently a draft token is rejected in the verification phase, triggering a rollback. To this end, we derive the probability that a token sampled from $q$ is rejected during verification, for a target distribution resulting from a deferral rule $r$.color=RoyalBlue!35!white]nit, in Section 3 we did not keep a separate section on “deferral risk for token-level cascades”. Hari: combined with next section!
@@ 4.3 Optimal speculative cascade deferral
For a given prefix $x_{<t}$, and target distribution $\pi=(1-r(x_{<t}))\cdot q_{t}+r(x_{<t})\cdot p_{t}$,color=green!30!white]$\pi_{t}$? the probability of a token drawn from draft distribution $q_{t}$ being rejected is equal to: $r(x_{<t})\cdot D_{\textup{{TV}}}(p_{t},q_{t}),$ where $D_{\textup{{TV}}}(p,q)=\sum_{v\in\mathscr{V}}\max\{0,p(v)-q(v)\}$ is the TV distance between $p$ and $q$.
Intuitively, whenever $r(x_{<t})=0$, $\pi(v)=q_{t}(v)$, and therefore there is no rejection or roll-back; when $r(x_{<t})=1$, the rejection rate equals $D_{\textup{{TV}}}(p_{t},q_{t})$.
@@
$\min_{r}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(v,q_{t})+r(x_{<t})\cdot\ell(v,p_{t})\big{)}\Big{]}\hskip 2.84544pt\text{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ r(x_{<t})\cdot D_{\textrm{TV}}(p_{t},q_{t})\,\leq\,B,$ (7)
for some budget $B>0$. Equivalently, one may minimize an unconstrained objective similar to equation 3, for suitable cost parameter $\alpha>0$ (see §C.4):
$L_{\rm spec}(r;x_{<t})$ $=\mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(v,q_{t})+r(x_{<t})\cdot\big{(}\ell(v,p_{t})+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})\big{)}\big{]},$ (8)
Contrasting equation 8color=green!30!white]nit: sometimes we use "equation 3". Sometimes we use (3). with the deferral risk in equation 3 for a sequential cascade, a key difference is that the cost of deferring to the larger model is no longer a constant, but depends on the similarity between $q_{t}$ and $p_{t}$, as measured by the TV distance between them.
color=RoyalBlue!35!white]nit, have we used the phrase “speculative deferral” in the text? what about “optimal speculative cascade deferral”?. Hari: done! We next derive the optimal deferral rule for equation 8, and construct a feasible estimator for it.
@@
The minimizer of equation 8 is of the form:
$r^{*}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(v,q_{t})\right]\,>\,\mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(v,p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (9)
@@
Plug-in estimator for equation 9. The optimal rule requires estimating expectations with respect the ground-truth distribution $\mathbb{P}(\cdot|x_{<t}).$ We employ similar plug-in estimators as the ones used with sequential cascades (§3). When $\ell=\ell_{\text{0-1}}$, we replace the expected 0-1 loss with (one minus) the maximum probability from the model, giving us:
$\textstyle{\hat{r}_{\rm\tt OPT}(x_{<t})=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).}$ (10)
@@
$L_{\rm spec}(\hat{r}_{\rm\tt OPT};x_{<t})-\min_{r}\,L_{\rm spec}(r;x_{<t})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,\max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$